# Personalized Speech Emotion Recognition in Human-Robot Interaction using Vision Transformers

### v1

@@ Abstract
Emotions are an essential element in verbal communication, so understanding individuals’ affect during a human-robot interaction (HRI) becomes imperative. This paper investigates the application of vision transformer models, namely ViT (Vision Transformers) and BEiT (BERT Pre-Training of Image Transformers) pipelines, for Speech Emotion Recognition (SER) in HRI. The focus is to generalize the SER models for individual speech characteristics by fine-tuning these models on benchmark datasets and exploiting ensemble methods. For this purpose, we collected audio data from different human subjects having pseudo-naturalistic conversations with the NAO robot. We then fine-tuned our ViT and BEiT-based models and tested these models on unseen speech samples from the participants. In the results, we show that fine-tuning vision transformers on benchmark datasets and and then using either these already fine-tuned models or ensembling ViT/BEiT models gets us the highest classification accuracies per individual when it comes to identifying four primary emotions from their speech: neutral, happy, sad, and angry, as compared to fine-tuning vanilla-ViTs or BEiTs.
@@ I Introduction
The increasing integration of social robots across various sectors, from healthcare to customer service, underscores their potential to revolutionize human-machine interactions [1, 2, 3, 4]. A crucial factor in their success lies in their ability to perceive and respond appropriately to human emotions, facilitating meaningful and engaging interactions [2, 5]. In this context, SER emerges as a critical field within human-computer interaction [6]. By enabling machines to understand and respond to the emotional nuances embedded in human speech (affective speech), SER can transform our interactions with technology, fostering more natural and empathetic communication [6]. When social robots can accurately interpret affective speech, they can adapt their behavior and responses, leading to more personalized and impactful human interactions [2]. This emotional connection ability holds tremendous potential for enhancing the effectiveness and acceptance of social robots in various real-world applications.
@@
Figure 1: The workflow showing speech emotion recognition during human-robot interaction.
@@
  • To the best of our knowledge, this work is among the first/earliest to investigate vision transformer-based models (both ViT and BEiT) for SER in the context of pseudo-naturalistic verbal HRi.
  • We show that personalization of SER models can be done by fine-tuning ViT and BEiT models on benchmark datasets and then fine-tuning on participant data and/or ensembling these vision transformer models.
  • Lastly, we also achieve SOTA performance on some benchmark datasets by a full fine-tuning of the vision transformer-based models.
This paper has been arranged in the following manner: Section II outlines the background literature supporting this work. Section III descibes the methodology, which includes the data acquisition (Section III-A), description about mel-spctrograms (Section III-B), datasets used (Section III-C), and problem formulation (Section III-D). This is followed by Section IV, which discusses the results we obtained, and followed by the conclusion in Section VI.
@@ II Related Works
The authors in [20] proposed to use CNN and RNN pipelines along with data augmentation techniques to improve the robustness of these models. This robustness was crucial for a human-robot interaction scenario with robot’s ego noise. Further, the authors in [21] propose a machine learning pipeline for SER. Their approach involves using personalized and non-personalized features for SER. However, neither of these papers contributes to evaluating transformer-based architectures, which are currently state-of-the-art (SOTA) in numerous fields of study [19].
@@ III-A Data Acquisition
Six neurotypical adults were recruited to participate in a human-robot interaction study to classify their speech into four primary emotions. This one-to-one HRI involves two-way communication between the robot and the participants. Each participant asks pre-defined questions as shown in Figure 1. The robot responds with appropriate answers to those questions and asks those questions back to the participant. The participants’ replies are not pre-defined. They were asked to reply to the robot’s questions with short answers. For each of these question-and-answer pairs, each participant was asked to speak in an emotional tone depicting one of the four primary emotions, i.e., neutral, happy, sad, and angry. The voices of the participants were recorded during this pseudo-natural human-robot interaction where the questions that the participant asks were pre-defined but their answers weren’t.
@@ III-B Mel Spectrogram
In this paper, since we are using vision based models, we convert the sound signals to 2D images. This is where we leverage the use of mel spectrograms. The mel spectrogram is used usually for better perception of sounds by humans. Considering $f$ as the normal frequency, the frequency on the mel scale ($m$) will be given by [31, 32, 33]:

### v2

@@ Abstract
Emotions are an essential element in human verbal communication, therefore it is important to understand individuals’ affect during human-robot interaction (HRI). This paper investigates the application of vision transformer models, namely ViT (Vision Transformers) and BEiT (Bidirectional Encoder Representations from Pre-Training of Image Transformers) pipelines for Speech Emotion Recognition (SER) in HRI. The focus is to generalize the SER models for individual speech characteristics by fine-tuning these models on benchmark datasets and exploiting ensemble methods. For this purpose, we collected audio data from several human subjects having pseudo-naturalistic conversations with the NAO social robot. We then fine-tuned our ViT and BEiT-based models and tested these models on unseen speech samples from the participants in order to dentify four primary emotions from speech: neutral, happy, sad, and angry. The results show that fine-tuning vision transformers on benchmark datasets and then using either these already fine-tuned models or ensembling ViT/BEiT models results in higher classification accuracies than fine-tuning vanilla-ViTs or BEiTs.
@@ I Introduction
The increasing integration of social robots across various sectors, from healthcare to customer service, underscores their potential to revolutionize human-machine interaction [1, 2, 3, 4]. A crucial factor in their application success is the ability to perceive and respond appropriately to human emotions, facilitating meaningful and engaging interactions [2, 5, 6, 7]. In this context, Speech Emotion Recognition (SER) emerges as a critical field within human-computer interaction [8]. By enabling machines to understand and respond to the emotional nuances embedded in human speech (affective speech), SER can transform our interactions with technology, fostering more natural and empathetic communication [8]. When social robots can accurately interpret affective speech, they can adapt their behavior and responses, leading to more personalized and impactful human interactions [2]. This emotional connection ability holds tremendous potential for enhancing the effectiveness and acceptance of social robots in various real-world applications.
@@
Figure 1: The experimental workflow showing our speech emotion recognition pipeline during human-robot interaction.
@@
  • This work is among the first to investigate vision transformer-based models (both ViT and BEiT) for SER in the context of pseudo-naturalistic verbal HRI.
  • We show that personalization of SER models can be done by fine-tuning ViT and BEiT models on benchmark datasets and then further fine-tuning these on participant data and through ensembling the models.
  • Lastly, we also achieve state-of-the-art (SOTA) performance on the RAVDESS and TESS datasets by a full fine-tuning of the vision transformer-based models.
This paper has been arranged in the following manner: Section II outlines the background literature supporting this work. Section III descibes the methodology, which includes the data acquisition (Section III-A), description about mel-spectrograms (Section III-B), datasets used (Section III-C), and problem formulation (Section III-D). This is followed by Section IV, which discusses the results we obtained, and followed by the conclusion in Section VI.
@@ II Related Works
The authors in [22] proposed to use CNN and RNN pipelines along with data augmentation techniques to improve the robustness of these models. This robustness was crucial for a human-robot interaction scenario with robot’s ego noise. Further, the authors in [23] propose a machine learning pipeline for SER. Their approach involves using personalized and non-personalized features for SER. However, neither of these papers contributes to evaluating transformer-based architectures, which are currently SOTA in numerous fields of study [21].
@@ III-A Data Acquisition
Six neurotypical adults were recruited to participate in a human-robot interaction study to classify their speech into four primary emotions. This one-to-one HRI involves two-way communication between the robot and the participants. Each participant asks pre-defined questions as shown in Figure 1. These questions had been used for our previous studies during HRI [33]. The following are the questions we asked the participants to ask the robot:
  • Hi. What’s your name?
  • How are you doing?
  • Did you do anything fun yesterday?
  • What do you like doing?
  • Any plans for the weekend?
The robot responds with appropriate answers to those questions and asks those questions back to the participant. The participants’ replies are not pre-defined. They were asked to reply to the robot’s questions with short answers. For each of these question-and-answer pairs, each participant was asked to speak in an emotional tone depicting one of the four primary emotions, i.e., neutral, happy, sad, and angry. The voices of the participants were recorded during this pseudo-natural human-robot interaction where the questions that the participant asks were pre-defined but their answers weren’t.
@@ III-B Mel Spectrogram
In this paper, since we are using vision based models, we convert the sound signals to 2D images. This is where we leverage the use of mel spectrograms. The mel spectrogram is used for better perception of sounds by humans. Considering $f$ as the normal frequency, the frequency on the mel scale ($m$) will be given by [34, 35, 36]:
@@ VI Conclusion
In the future, we would like to recruit more human participants and collect data across different populations, including both neurotypical and neurodivergent populations. We would also like to examine multiple data modalities and examine how speech emotion correlates to modalities such as facial videos and physiological signals.